# Feasibility of Learning Weighted Automata on a Semiring

### v3

@@ 5. Conclusion and open questions
We have pictured a landscape for learning weighted automata on any semiring and shown that an approach à la Angluin can only work in general on a restricted class of functions. This paper opens many questions, both general and specific. First, for functions that are not weakly guessable, one can ask whether there is still a way to find appropriate states (not from rows of the Hankel matrix) by strengthening the queries. Next, for weakly guessable functions, one can ask whether there is a way to design strategies to update $Q$ and $T$ (other than closure and adding suffixes of counter-examples to $T$) that would ensure to find a suitable hypothesis automaton. In particular, is there a way to choose “cleverly” the coefficients of the linear combinations under consideration? For example, the algorithm of [21] can fail for functions that are not guessable (since there is no condition on how the coefficients are chosen). For guessable functions, a naive way to ensure termination is to make sure to enumerate all the words in $T$. But is there a better way? And if so, it would be interesting to see if the class of functions learnable by a polynomial time algorithm can be characterised. Regarding the general hierarchy of functions, there are several classes for which we neither have an example of non-emptiness or an argument for collapse. It would also be of interest to understand how termination of the algorithm from [21] relates to strong guessability.
@@ Acknowledgment
This work has been supported by the EPSRC grant EP/T018313/1 and a Turing-Manchester Exchange Fellowship, The Alan Turing Institute. We also thank Nathanaël Fijalkow for useful discussions on the topic.
@@ Appendix A Dual approach using columns
In this section, we present (for clarity) the natural dual definitions to those provided in Section 2 and Section 3, working with the transpose of the Hankel matrix. Thus, we work with (partial) columns instead of (partial) rows and consider right multiplication instead of left multiplication. As a result, the role of final and initial states is swapped in our constructions. $S$ still denotes a semiring and $\Sigma$ an alphabet.
@@
Given a set $Y$, an element $x$ of $S^{Y}$ and $\lambda$ in $S$, $x\otimes\lambda$ denotes the element of $S^{Y}$ computed from $x$ by multiplying every component of $x$ by $\lambda$ on the right. Given a set $X=\{x_{1},x_{2},\ldots\}$ of elements of $S^{Y}$, a right-linear combination over $X$ is one of the form $(x_{1}\otimes\lambda_{1})\oplus(x_{2}\otimes\lambda_{2})\oplus\ldots$ for some $\lambda_{1},\lambda_{2},\ldots$ in $S$, such that only a finite number are different from $0_{S}$. The right-semimodule generated by $X$ is defined as all the elements of $S^{Y}$ that can be written as a right-linear combination over $X$. An element of the right-semimodule generated by $X$ will simply be said to be right-generated by $X$.
We will use the notation $[x]$ for elements of $S^{\Sigma^{*}}$; thus for a word $w\in\Sigma^{*}$ we write $[x]_{w}$ to denote the $w$-th component of $[x]$. The restriction of an element of $S^{\Sigma^{*}}$ to entries indexed by a subset $Z\subseteq\Sigma^{*}$ will be denoted by $[x]_{Z}$. Given an element $[x]\in S^{\Sigma^{*}}$, and a letter $a$ of $\Sigma$, we denote by $[ax]$ the element of $S^{\Sigma^{*}}$ defined by $[ax]_{w}=[x]_{wa}$ for all words $w$. If $x$ is a word and $f$ and $F$ are clear from context, we shall write $[x]$ to denote the element of $S^{\Sigma^{*}}$ indexed by $x$ in $F$. Note that if $x$ is a word, $f$ a function and $F$ its Hankel matrix, this definition is consistent: $[ax]_{u}=F_{u,ax}=f(uax)=F_{ua,x}=[x]_{ua}$. For a subset $X\subseteq S^{\Sigma^{*}}$ and $[x]$ in $X$, we will sometimes simplify notation by writing $x\in X$.
Let $Z\subseteq\Sigma^{*}$ and $f:\Sigma^{*}\rightarrow S$. A subset $X\subseteq S^{\Sigma^{*}}$ is said to be: (i) right-closed if for all $[x]$ in $X$ and $a\in\Sigma$, $[ax]$ is right-generated by $X$; (ii) right-closed on $Z$ if for all $[x]$ in $X$ and $a\in\Sigma$, $[ax]_{Z}$ is right-generated by the elements of $X$ restricted to $Z$; (iii) column-generating for $f$ if all the columns of the Hankel matrix of $f$ are right-generated by $X$. By an abuse of notation, we say that a set of words $Q$ is right-closed (resp. right-closed on $Z$, column-generating for $f$) if the set $X=\{[q]:q\in Q\}$ has this property.
@@
Given a function $f:\Sigma^{*}\rightarrow S$, a finite subset $T$ of $S^{\Sigma^{*}}$ and a subset $Q$ of $\Sigma^{*}$, we define $\Gamma_{Q,T}\subseteq S^{T}\times S^{T\times\Sigma\times T}$ to be the set of tuples $\lambda=(\lambda_{q}^{\varepsilon},\lambda_{q,a,p})_{q\in T,a\in\Sigma,p\in T}$ such that:
  • $[\varepsilon]_{Q}=\bigoplus_{q\in T}[q]_{Q}\otimes\lambda_{q}^{\varepsilon}$, and
  • for all $q\in T$ and all $a\in\Sigma$, $[aq]_{Q}=\bigoplus_{p\in T}[p]_{Q}\otimes\lambda_{q,a,p}$.
Note that $\Gamma_{Q,T}$ is non-empty if and only if $T$ is right-closed on $Q$ and $[\varepsilon]_{Q}$ is right-generated by $T$ restricted to $Q$. Note also that if $Q\subseteq Q’$ ’, then $\Gamma_{Q,T}\supseteq\Gamma_{Q^{\prime},T}$. If $Q=\Sigma^{*}$, we simply write $\Gamma_{T}$.
@@
Given a finite subset $T$ of $S^{\Sigma^{*}}$ and a subset $Q$ of $\Sigma^{*}$, for each $\lambda\in\Gamma_{Q,T}$, we define a finite weighted automaton $\mathcal{H}^{co}_{Q,T,\lambda}$ with:
@@
  • initial-state vector $([q]_{\varepsilon})_{q\in T}$;
  • final-state vector $(\lambda^{\varepsilon}_{q})_{q\in T}$; and
  • for all $q,p\in T$ and all $a\in\Sigma$, a transition from $p$ to $q$ labelled by $a$ with weight $\lambda_{q,a,p}$.
If $Q=\Sigma^{*}$, we say $\mathcal{H}^{co}_{Q,T,\lambda}$ is a co-Hankel automaton and write simply $\mathcal{H}^{co}_{T,\lambda}$ to reduce notation. If $T$ is a finite set of columns of the Hankel matrix and $Q$ is finite, we say that $\mathcal{H}^{co}_{Q,T,\lambda}$ is a co-hypothesis automaton, and view $T$ as a finite set of words by identifying with an appropriate index set.
@@
Let $f:\Sigma^{*}\rightarrow S$ be a function. There exists a finite set $T\subseteq S^{\Sigma^{*}}$ that is right-closed and right-generates $\langle\varepsilon\rangle$ if and only if $f$ is computed by a finite weighted automaton $\mathcal{A}$ over $S$. Specifically, if there exists a finite set $T\subseteq S^{\Sigma^{*}}$ that is right-closed and right-generates $\langle\varepsilon\rangle$, then $\mathcal{H}^{co}_{T,\lambda}$ computes $f$ for all $\lambda\in\Gamma_{T}\neq\emptyset$, whilst if $f$ is computed by a finite weighted automaton $\mathcal{A}$ over $S$, then the set of all $[q]$ where $q$ is a state of $\mathcal{A}$ and $[q]_{w}$ is the weight of $w$ to $q$ in $\mathcal{A}$ is right-closed and right-generates every column of the Hankel matrix of $f$.
@@
A function $f:\Sigma^{*}\rightarrow S$ is said to be weakly co-guessable if there exist a finite set $T$ of columns of the Hankel matrix of $f$ such that $\Gamma_{T}$ is non-empty.
@@
A function $f:\Sigma^{*}\rightarrow S$ is co-guessable if there exist finite subsets $Q,T\subseteq\Sigma^{*}$ such that $\Gamma_{Q,T}=\Gamma_{T}\neq\emptyset$. We say that the pair $(Q,T)$ witnesses that $f$ is co-guessable. Furthermore, we say that a co-guessable function $f$ is strongly co-guessable if for all finite sets $T\subseteq\Sigma^{*}$ with $\Gamma_{T}\neq\emptyset$ there exists a finite set $Q\subseteq\Sigma^{*}$ such that $(Q,T)$ witnesses that $f$ is co-guessable.
@@
Let $T$ be a finite subset of $\Sigma^{*}$. A total $T$-chain is an infinite descending chain of solution spaces $\Gamma_{Q_{0},T}\supsetneq\Gamma_{Q_{1},T}\supsetneq\Gamma_{Q_{2},T}\supsetneq\cdots$ with $Q_{i}\subseteq Q_{i+1}$ for all $i\geq 0$ and such that each word in $\Sigma^{*}$ is contained in some $Q_{i}$. We say that a function $f:\Sigma^{*}\rightarrow S$ is column-bound if there exists a finite set of words $T$ such that $\Gamma_{T}\neq\emptyset$ and there is no total $T-$chain. We say that $f:\Sigma^{*}\rightarrow S$ is strongly column-bound if for every finite set of words $T$ such that $\Gamma_{T}\neq\emptyset$ there is no total $T$-chain.
@@ Appendix B Examples
Let $S$ be one of $\mathbb{N}_{\max}$, $\mathbb{Z}_{\max}$, or $\mathbb{R}_{\max}$, and consider the function $f_{1}$ computed by the automaton depicted in Figure 1(b) and $F$ its Hankel matrix over $S$. Let $w=ua^{n}$ and $w^{\prime}=a^{n^{\prime}}v$ where $u,v\in\Sigma^{*}$ are such that $u$ does not end with an $a$ and $v$ does not start with an $a$. The corresponding entry of the Hankel matrix is given by $F_{w,w^{\prime}}=f(ww^{\prime})={\max}\{f_{1}(u),n+n^{\prime},f_{1}(v)\}$. We claim that $f_{1}$ is neither weakly guessable nor weakly co-guessable. Suppose for contradiction that $f_{1}$ is weakly guessable. By Theorem 3.5 there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the rows of the Hankel matrix indexed by $W$ are row-generating. Let us write $w_{i}=u_{i}a^{n_{i}}$ for $i=1,\ldots,k$, where $u_{i}$ does not end with an $a$ and $n_{i}\geq 0$, and choose a positive integer $N$ such that $N\geq{\max}_{i\in\{1,\ldots,k\}}\{f_{1}(u_{i}),n_{i}\}$. Consider the row of the Hankel matrix indexed by $w=a^{N+1}b$. Since the rows indexed by $W$ are assumed to be row-generating, we have: $\langle w\rangle=\bigoplus_{i=1}^{k}\lambda_{i}\otimes\langle w_{i}\rangle$ for some $\lambda_{i}\in S$. Now set $u=ba^{N}$ and $v=ba^{N+1}$. We have:
@@
Let $S$ be one of $\mathbb{N}$, $\mathbb{R}_{\geq 0}$ and consider the function $f^{\prime}_{1}$ computed by the automaton depicted in Figure 1(d) and $F$ its Hankel matrix. Suppose that $f^{\prime}_{1}$ is weakly guessable. Then there exist a finite set of words $Q=\{w_{1},w_{2},\ldots,w_{k}\}$ such that the rows indexed by $Q$ left-generate the rows of $F$. Without loss of generality, we may assume that $w_{1}=\varepsilon$ and that $w_{i}$ is a power of $a$ for $i=1,\ldots m\leq k$. Let $w=a^{n}$ for some positive integer $n$. Then by assumption, considering columns $\varepsilon$ and $a$ of $F$, there exist $\lambda_{1},\ldots,\lambda_{k}$ in $S$ such that:
@@
For $S=\mathcal{P}_{\rm fin}(\Sigma^{*})$, consider the function $f^{\prime\prime}_{1}$ computed by the automaton depicted in Figure 1(f) and $F$ its Hankel matrix. If a row $\langle w\rangle$ of $F$ is left-generated by a finite set of rows, this in particular means (by restricting attention to column $\varepsilon$) that $\{a^{|w|_{a}},b^{|w|_{b}}\}$ can be written as $\bigcup_{i=1}^{k}X_{i}\{a^{n_{i}},b^{m_{i}}\}$ for some finite sets $X_{i}$ and non-negative integers $n_{i},m_{i}$. This is only possibly if exactly one of the sets $X_{i}$ is non-empty, and for this value $i$ we have that $X_{i}=\{\varepsilon\}$ and $n_{i}=|w|_{a}$ and $m_{i}=|w|_{b}$. It follows from this that for all pairs of non-negative integers $(n,m)$ a row-generating set of words must contain at least one word with $|w|_{a}=n$ and $|w|_{b}=m$.
@@
Let $S=\mathbb{Z}_{\max}$ or $S=\mathbb{R}_{\max}$ and consider the function $f_{2}$ computed by the automaton depicted in Figure 1(a), but viewed as an automaton over $S$ and let $F$ denote its Hankel matrix.

### v4

@@ 5. Conclusion and open questions
We have pictured a landscape for learning weighted automata on any semiring and shown that an approach à la Angluin can only work in general on a restricted class of functions. This paper opens many questions, both general and specific. First, for functions that are not weakly guessable, one can ask whether there is still a way to find appropriate states (not from rows of the Hankel matrix) by strengthening the queries. Next, for weakly guessable functions, one can ask whether there is a way to design strategies to update $Q$ and $T$ (other than closure and adding suffixes of counter-examples to $T$) that would ensure to find a suitable hypothesis automaton. In particular, is there a way to choose “cleverly” the coefficients of the linear combinations under consideration? For example, the algorithm of [36] can fail for functions that are not guessable (i.e. if for some finite set $Q$ at every Test step the counter example provided by the teacher allows for the possibility that the closure strategy produces a solution that is not in $\Lambda_{Q}$). For guessable functions, a naive way construct an algorithm that terminates with success is to use the update strategy suggested by the proof of Proposition 3.18. But is there a better way? And if so, it would be interesting to see if the class of functions learnable by a polynomial time algorithm can be characterised. Regarding the general hierarchy of functions, there is a class for which we neither have an example of non-emptiness or an argument for collapse: the functions that are guessable but not strongly guessable (and its dual). Finally, it would be interesting to study the notion of uniform strong guessability introduced in Remark 4.2 and in particular, try to characterise the class of functions lying in the intersection of the six classes we have introduced. It would also be of interest to understand how termination of the algorithm from [36] relates to strong guessability.
@@ Acknowledgment
This work has been supported by the EPSRC grant EP/T018313/1 and a Turing-Manchester Exchange Fellowship, The Alan Turing Institute. We also thank Nathanaël Fijalkow for useful discussions on the topic, and the anonymous reviewers for their deep and useful comments.
@@ Appendix A Dual approach using columns
In this section, we present (for clarity) the natural dual definitions to those provided in Section 2 and Section 3 working with (partial) columns instead of (partial) rows and consider right multiplication instead of left multiplication. As a result, the role of final and initial states is swapped in our constructions. $S$ still denotes a semiring and $\Sigma$ an alphabet.
@@
Given a set $Y$, $S^{Y}$ is a right $S$-module via the following action: for each element $\underline{x}\in S^{Y}$ and $\lambda$ in $S$, $\underline{x}\otimes\lambda$ denotes the element of $S^{Y}$ computed from $\underline{x}$ by multiplying every component of $\underline{x}$ by $\lambda$ on the right. Given a subset $X\subseteq S^{Y}$, a right-linear combination over $X$ is one of the form $\bigoplus_{\underline{x}\in X}(\underline{x}\otimes\lambda_{x})$ for some $\lambda_{\underline{x}}\in S$, with only finitely many $\lambda_{\underline{x}}$ different from $0_{S}$. The right-semimodule generated by $X$ is defined as all the elements of $S^{Y}$ that can be written as a right-linear combination over $X$. An element of the right-semimodule generated by $X$ will simply be said to be right-generated by $X$.
As before, we will use the notation $\underline{x}$ for elements of $S^{\Sigma^{*}}$ and $\underline{x}_{Z}$ for the restriction of $\underline{x}$ to entries indexed by a subset $Z\subseteq\Sigma^{*}$, remembering that the action of $S$ is now on the right. We also define a left action of $\Sigma^{*}$ on $S^{\Sigma^{*}}$ as follows: given an element $\underline{x}\in S^{\Sigma^{*}}$, and $u\in\Sigma^{*}$, we denote by $u\cdot\underline{x}$ the element of $S^{\Sigma^{*}}$ defined by $(u\cdot\underline{x})_{w}=\underline{x}_{wu}$ for all words $w$. It will also be convenient to have a notation for elements of $S^{\Sigma^{*}}$ that arise as columns of a given Hankel matrix $F$. If $v$ is a word, we shall write $[v]$ to denote the infinite column indexed by $v$ in $F$, viewed as an element of $S^{\Sigma^{*}}$. Note that the left action of $\Sigma^{*}$ on $[v]$ corresponds to the obvious left action of $\Sigma^{*}$ on $v$, that is: $(u\cdot[v])_{w}=[v]_{wu}=F_{wu,v}=f(wuv)=F_{w,uv}=[uv]_{w}$.
Let $Z\subseteq\Sigma^{*}$ and $f:\Sigma^{*}\rightarrow S$. A subset $X\subseteq S^{\Sigma^{*}}$ is said to be: (i) right-closed if for all $\underline{x}$ in $X$ and $a\in\Sigma$, $a\cdot\underline{x}$ is right-generated by $X$; (ii) right-closed on $Z$ if for all $\underline{x}$ in $X$ and $a\in\Sigma$, $(a\cdot\underline{x})_{Z}$ is right-generated by the elements of $X$ restricted to $Z$; (iii) column-generating for $f$ if all the columns of the Hankel matrix of $f$ are right-generated by $X$. By an abuse of language, we say that a set of words $T$ is right-closed (resp. right-closed on $Z$, column-generating for $f$) if the set $\underline{T}=\{[t]:t\in T\}$ has this property.
@@
Given a function $f:\Sigma^{*}\rightarrow S$, a finite non-empty subset $T$ of $S^{\Sigma^{*}}$ and a non-empty subset $Q$ of $\Sigma^{*}$, we define the right-solution set $\Gamma_{Q,T}$ to be set of elements $\underline{\gamma}\in S^{T\cup(T\times\Sigma\times T)}$ satisfying:
  • $[\varepsilon]_{Q}=\bigoplus_{\underline{t}\in T}\underline{t}_{Q}\otimes\gamma_{\underline{t}}$, and
  • for all $\underline{t}\in T$ and all $a\in\Sigma$, $(a\cdot\underline{t})_{Q}=\bigoplus_{\underline{s}\in T}\underline{s}_{Q}\otimes\underline{\gamma}_{\underline{s},a,\underline{t}}$.
If $Q=\Sigma^{*}$, we will simply write $\Gamma_{T}$ instead of $\Gamma_{Q,T}$. Moreover, if $R\subseteq\Sigma^{*}$, then to reduce notation we will also frequently write simply $\Gamma_{Q,R}$ and $\Gamma_{R}$ instead of $\Gamma_{Q,T}$ and $\Gamma_{T}$ where $T=\{[t]:t\in R\}$.
It is clear from the definitions above that $\Gamma_{Q,T}$ is non-empty if and only if $[\varepsilon]_{Q}$ is right-generated by $T$ restricted to $Q$ and $T$ is right-closed on $Q$. Also, if $Q^{\prime}\subseteq Q$, then $\Gamma_{Q^{\prime},T}\supseteq\Gamma_{Q,T}$. However, if $T^{\prime}$ is a finite set contained in or contained by $T$, there is no containment between $\Gamma_{Q,T}$ (which lies in $S^{T\cup(T\times\Sigma\times T)}$) and $\Gamma_{Q,T^{\prime}}$ (which lies in $S^{T^{\prime}\cup(T^{\prime}\times\Sigma\times T^{\prime})}$).
@@
Let $f:\Sigma^{*}\rightarrow S$ be a function and suppose that $T$ is a finite subset of $S^{\Sigma^{*}}$ and $Q$ is a subset of $\Sigma^{*}$ such that $\Gamma_{Q,T}\neq\emptyset$. Then for each $\gamma\in\Gamma_{Q,T}$, we define a finite weighted automaton $\mathcal{H}^{co}_{Q,T,\underline{\gamma}}$ with:
@@
  • initial-state vector $(\underline{t}_{\varepsilon})_{\underline{t}\in T}$;
  • final-state vector $\underline{\gamma}_{T}$; and
  • for all $\underline{s},\underline{t}\in T$ and all $a\in\Sigma$, a transition from state $\underline{s}$ to state $\underline{t}$ labelled by the letter $a$ with weight $\underline{\gamma}_{\underline{s},a,\underline{t}}$.
If $Q=\Sigma^{*}$, we say $\mathcal{H}^{co}_{Q,T,\underline{\gamma}}$ is a co-Hankel automaton and write simply $\mathcal{H}^{co}_{T,\underline{\gamma}}$ to reduce notation. If $T$ is a finite set of columns of the Hankel matrix and $Q$ is finite, we say that $\mathcal{H}^{co}_{Q,T,\underline{\gamma}}$ is a co-hypothesis automaton; in this case we may view $T$ as a finite set of words by identifying with an appropriate index set note: there need not be a unique set of words.
@@
Let $f:\Sigma^{*}\rightarrow S$ be a function. The following are equivalent:
  1. (1)
    there exists a finite subset $T\subseteq S^{\Sigma^{*}}$ that is both right-closed and right-generates the column indexed by $\varepsilon$ in the Hankel matrix of $f$ (i.e. $\Gamma_{T}\neq\emptyset$ for some finite subset $T\subseteq S^{\Sigma^{*}}$);
  2. (2)
    $f$ is computed by a finite-state weighted automaton $\mathcal{A}$ over $S$.
Specifically, if $\Gamma_{T}\neq\emptyset$, then $\mathcal{H}^{\rm co}_{T,\underline{\gamma}}$ computes $f$ for all $\underline{\gamma}\in\Gamma_{T}$, whilst if $f$ is computed by the finite-state weighted automaton $\mathcal{A}$ over $S$ with state set $T$, then $\Gamma_{\underline{T}}\neq\emptyset$ for $\underline{T}=\{\underline{t}:t\in T\}\subseteq S^{\Sigma^{*}}$ where for each $w\in\Sigma^{*}$ we define $\underline{t}_{w}$ to be the value on input $w$ computed by the automaton $\mathcal{A}_{t}$ obtained from $\mathcal{A}$ by making $t$ the unique final state with weight $1_{S}$ and all other weights are the same).
@@
A function $f:\Sigma^{*}\rightarrow S$ is said to be weakly co-guessable if there exists a finite set $T$ of columns of the Hankel matrix of $f$ such that $\Gamma_{T}$ is non-empty.
@@
A function $f:\Sigma^{*}\rightarrow S$ is co-guessable if there exist finite subsets $Q,T\subseteq\Sigma^{*}$ such that $\Gamma_{Q,T}=\Gamma_{T}\neq\emptyset$. We say that the pair $(Q,T)$ witnesses that $f$ is co-guessable. Furthermore, we say that a co-guessable function $f$ is strongly co-guessable if for all finite sets $T\subseteq\Sigma^{*}$, there exists a finite set $Q\subseteq\Sigma^{*}$ such that $\Gamma_{Q,T}=\Gamma_{T}$.
@@
Let $T$ be a finite subset of $\Sigma^{*}$. A total $T$-chain is an infinite descending chain of right-solution sets $\Gamma_{Q_{0},T}\supsetneq\Gamma_{Q_{1},T}\supsetneq\Gamma_{Q_{2},T}\supsetneq\cdots$ with $Q_{i}\subseteq Q_{i+1}$ for all $i\geq 0$ and such that each word in $\Sigma^{*}$ is contained in some $Q_{i}$. We say that a function $f:\Sigma^{*}\rightarrow S$ is column-bound if there exists a finite set of words $T$ such that $\Gamma_{T}\neq\emptyset$ and there is no total $T-$chain. We say that $f:\Sigma^{*}\rightarrow S$ is strongly column-bound if for every finite set of words $T$, there is no total $T$-chain.
@@ Appendix B Examples
Let $S$ be one of $\mathbb{N}_{\max}$, $\mathbb{Z}_{\max}$, or $\mathbb{R}_{\max}$, and consider the function $f_{1}$ computed by the automaton depicted in Figure 1(b) and $F$ its Hankel matrix over $S$. Thus $f_{1}(w)$ is equal to the length of the longest block of consecutives $a$’s in $w$. Let $w=ua^{n}$ and $w^{\prime}=a^{n^{\prime}}v$ where $u,v\in\Sigma^{*}$ are such that $u$ does not end with an $a$ and $v$ does not start with an $a$. The corresponding entry of the Hankel matrix is given by $F_{w,w^{\prime}}=f(ww^{\prime})={\max}\{f_{1}(u),n+n^{\prime},f_{1}(v)\}$. We claim that $f_{1}$ is neither weakly guessable nor weakly co-guessable. Suppose for contradiction that $f_{1}$ is weakly guessable. By Theorem 3.6 there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the rows of the Hankel matrix indexed by $W$ are row-generating. Let us write $w_{i}=u_{i}a^{n_{i}}$ for $i=1,\ldots,k$, where $u_{i}$ does not end with an $a$ and $n_{i}\geq 0$, and choose a positive integer $N$ such that $N\geq{\max}_{i\in\{1,\ldots,k\}}\{f_{1}(u_{i}),n_{i}\}$. Consider the row of the Hankel matrix indexed by $w=a^{N+1}b$. Since the rows indexed by $W$ are assumed to be row-generating, we have: $\langle w\rangle=\bigoplus_{i=1}^{k}\lambda_{i}\otimes\langle w_{i}\rangle$ for some $\lambda_{i}\in S$. Now set $u=ba^{N}$ and $v=ba^{N+1}$. We have:
@@
Let $S$ be one of $\mathbb{N}$, $\mathbb{R}_{\geq 0}$ and consider the function $f^{\prime}_{1}$ computed by the automaton depicted in Figure 1(d) and $F$ its Hankel matrix. Thus $f^{\prime}_{1}(w)$ is equal $12^{n}-$ 1 if $w=a^{n}$ for some $n>0$ and 00 otherwise. Suppose that $f^{\prime}_{1}$ is weakly guessable. Then there exist a finite set of words $Q=\{w_{1},w_{2},\ldots,w_{k}\}$ such that the rows indexed by $Q$ left-generate the rows of $F$. Without loss of generality, we may assume that $w_{1}=\varepsilon$ and that $w_{i}$ is a power of $a$ for $i=1,\ldots m\leq k$. Let $w=a^{n}$ for some positive integer $n$. Then by assumption, considering columns $\varepsilon$ and $a$ of $F$ and restricting attention to row $w$, there exist $\lambda_{1},\ldots,\lambda_{k}$ in $S$ such that:
@@
For $S=\mathcal{P}_{\rm fin}(\Sigma^{*})$, consider the function $f^{\prime\prime}_{1}$ computed by the automaton depicted in Figure 1(f) and $F$ its Hankel matrix. Thus $f^{\prime\prime}_{1}(w)=\{a^{|w|_{a}},b^{|w|_{b}}\}$. If a row $\langle w\rangle$ of $F$ is left-generated by a finite set of rows, this in particular means (by restricting attention to column $\varepsilon$) that $\{a^{|w|_{a}},b^{|w|_{b}}\}$ can be written as $\bigcup_{i=1}^{k}X_{i}\{a^{n_{i}},b^{m_{i}}\}$ for some finite sets $X_{i}$ and non-negative integers $n_{i},m_{i}$. This is only possibly if exactly one of the sets $X_{i}$ is non-empty, and for this value $i$ we have that $X_{i}=\{\varepsilon\}$ and $n_{i}=|w|_{a}$ and $m_{i}=|w|_{b}$. It follows from this that for all pairs of non-negative integers $(n,m)$ a row-generating set of words must contain at least one word with $|w|_{a}=n$ and $|w|_{b}=m$.
@@
Let $S=\mathbb{Z}_{\max}$ or $S=\mathbb{R}_{\max}$ and consider the function $f_{2}$ computed by the automaton depicted in Figure 1(a), but viewed as an automaton over $S$ and let $F$ denote its Hankel matrix. Thus $f_{2}(w)=|w|_{a}$ if $w$ starts with an $a$, $f_{2}(w)=|w|_{b}$ if $w$ starts with an $b$, and $f_{2}(\varepsilon)=-\infty$. Recalling that multiplication over this semiring is given by usual addition, if $w$ starts with $a$ we have $\langle w\rangle=(|w|_{a}-1)\otimes\langle a\rangle$. Likewise, if $w$ starts with $b$ we have $\langle w\rangle=(|w|_{b}-1)\otimes\langle b\rangle$. It follows from these simple observations that every row of the Hankel matrix is a scalar multiple of one of the rows $\langle\varepsilon\rangle,\langle a\rangle$ or $\langle b\rangle$. Moreover, it is not too difficult to see that none of these three can be expressed as a linear combination of the other two: for example the fact that $f_{2}(\varepsilon)=-\infty$ and $f_{2}(u)\neq-\infty$ for all non-empty words $u$ quickly demonstrates that $\langle\varepsilon\rangle$ cannot be generated by any combination of the other rows.